OpenStudy (anonymous):

-1/3k+-7/4=-9-3/5k k= thanks in advance

hero (hero):

Hi

OpenStudy (anonymous):

hi

hero (hero):

Is it: \[-\frac{1}{3}k + -\frac{7}{4} = -9 - \frac{3}{5}k\]

OpenStudy (anonymous):

yes

hero (hero):

Okay

hero (hero):

I will post a solution in a minute

hero (hero):

k = -435/16

OpenStudy (anonymous):

yes!!! Thank you